Suppose the short run Phillips Curve is given by:

Inflation = Expected Inflation + 0.2 - 4*Unemployment Rate

Assume that initially, people expect zero inflation.

a)Draw the short run Phillips Curve and the long run Phillips Curve on a graph

b)On the graph, represent what would happen in the short run if the government decided to run 4% inflation (setting inflation =0.04).

c)On the graph, represent what would happen in the long run if the government decided to run 4% inflation.
Answer #1

Long-Run Phillips Curve Inflation Short-Run Phillips Curve Unemployment The Phillips Curveb.)Inflation = Expected Inflation + 0.2-4* Unemployment rate.

Here, expected inflation =0, so, 4=0.2-4*Unemployment rate. Unemployment rate =-3.8/4=-0.95. Hence unemployment decreases at rate of 0.95 person.

c) Philips curve don't work in long run. This can be seen by stagflation in 1970. There will be no change in unemployment.